# The Impact of Train-Test Leakage on Machine Learning-based Android Malware Detection

### v1

@@ 5 Case Study 2: Graph Convolutional Networks for Android Malware Detection
5.1 Summary of the Research and Dataset Used
GCNs leverage graph structures as input, which allows for potentially capturing richer static code semantics, such as that expressed in an Android app’s ICFG. We use an Android app’s ICFG555 To obtain a single ICFG for an Android app, we combine the ICFG of each component of the app based on the life cycle semantics of how the methods of various components are called by the Android runtime. as the GCN model’s input, where each node represents an API call, with additional attributes including class path, method name, parameter type, and node degree. Following the standard GCN training process, we obtain a graph embedding for the ICFG, achieved through multiple rounds of message passing within the GCN, where each node’s embedding is updated based on its neighbors. Finally, a readout layer aggregates these node embeddings into a unified graph embedding for the entire ICFG. This component of the model can be seen as a graph encoder. The unified graph embedding learned by the encoder is then fed as input to the classification layer of the model. More details about the GCN model can be found in the Appendix.
We compare the GCN approach to a standard ML approach – random forest (RF). The RF model takes as input a feature vector representing an app. We extract DREBIN-like static features and rank them in decreasing order of their mutual information with the class label. We then select 2,970666The mutual information is close to 0 for features ranked lower. most informative features for training an RF model, which include permissions, intent actions, discriminative APIs, obfuscation signatures, and native code signatures.
@@ 5.1 Summary of the Research and Dataset Used
We collected Google Play apps released between January 2018 and December 2021 through AndroZoo [27], and used VirusTotal [29] results to label the apps. Using apps from a single market prevents the sampling bias [9] as mentioned in Section 2.2. We follow a sliding window approach to construct training, validation, and testing datasets, as shown in Figure 3. We first split malicious app data into batches along app release date, each batch initially consisting of 240 malicious apps. Then, we randomly sample benign apps from all the benign apps in a batch’s time frame to achieve a $6\%$ malware ratio. Specifically, for each 240 malicious apps batch, we sample 3,760 benign apps within the same time frame, for a total of 4,000 apps in a batch. This sampling based on release date avoids the spatial bias [10] as discussed in Section 2.2. We subsequently form windows of ten consecutive batches, with the first six batches used for training, the next two for validation, and the last two for testing. The complete dataset encompasses 24 batches, totaling 5,760 malicious and 90,240 benign apps, all chronologically ordered by their release date on Google Play. This dataset provides eight sliding windows (the window slides two batches each time) for evaluating machine learning models through different training, validation, and testing data. This sliding window approach, as the sampling approach for batch construction, also avoids the temporal bias [10] as discussed in Section 2.2.
5.2 Identification of Train-Test Leakage
@@
The two machine learning models our research compares, GCN and RF, use different input representations for an Android app. In GCN, an input app is represented as an ICFG adjacency matrix, together with a list of attributes for each node, whereas in RF an input app is represented as a feature vector. In prior discussion, we have explained the definition of train-test leakage when a feature vector is used as an app’s representation. These are apps in the testing data that share identical feature vectors with at least one app in the training data. We denote these leakage apps as $Leak_{\textit{fv}}$. While this definition works well for RF, it is not suitable to use in identifying train-test leakage for GCN-based models. Moreover, it is a non-trivial problem to determine whether two Android apps’ ICFGs are identical. However, as mentioned above, the GCN model consists of an encoder and a classification layer. The encoder generates a graph embedding, i.e., a vector of floating numbers, which is provided as input to the classification layer. Instead of comparing the ICFGs directly, we have chosen to use the computed graph embeddings of the ICFGs to identify train-test leakage. Specifically, given the embeddings $emb_{1}$ and $emb_{2}$ of two Android apps’ ICFGs, we compute cosine similarity between the two embeddings as
cosine $\textit{cosine similarity}=\frac{emb_{1}\cdot emb_{2}}{\|emb_{1}\|\|emb_{2}\|}$
@@
where $\cdot$ is the inner product operation and $\|v\|$ represents the length of a vector $v$. The cosine similarity measures how similar the two vectors are and ranges from 00 to $1$, where $1$ means the two vectors are identical, while 00 means that the two vectors are orthogonal and share no similarity.
@@
Note that due to GCN model’s stocasticity and the numerical precision limitations, even if we feed the exact same ICFG multiple times to a GCN, the model will not generally produce the exact same graph embeddings. Thus, for the purpose of identifying train-test leakage, we chose a threshold $M$. A test app is considered a leakage if there is at least one training app whose cosine similarity with the testing app is at least $M$. We denote these apps as $Leak_{\textit{emb}}$. For a given testing set, we choose $M$ so that the two sets of leakage apps in the testing data, $Leak_{\textit{fv}}$ and $Leak_{\textit{emb}}$, are most aligned. The $M$ value varies from one testing window (Figure 3) to another, but they are all very close to $1$ (e.g., $0.9998$). More details on how the threshold $M$ is determined can be found in Appendix.
@@
Figure 4 illustrates the ratio of leakage apps ($Leak_{\textit{fv}}$, $Leak_{\textit{emb}}$, and their union) on the testing data used in our experiments. A significant portion of apps in the testing data (oftentimes exceeding half of the total apps) are identified as train-test leakage. Since the two models give rise to two different sets of leakage apps, we need to remove the union of the two leakage sets from the testing data in order to compare the two models’ performance after leakage removal, just like what we did in Section 4.
5.3 Impact of the Leakage on Performance Results
Like in Section 4, we compare the performance measurements of the two models both before and after leakage apps are removed from testing data. Leakage portion is defined as $Leak_{\textit{fv}}\cup Leak_{\textit{emb}}$. Table III shows the results. For each testing window, the color-coded performance differences are included in parentheses, with green in favor of GCN and red in favor of RF. If we measure the performance on all the test data, RF consistently outperforms GCN based on the F1 score and balanced accuracy metric. However, once the leakage is removed, the comparative result flips: GCN outperforms RF in seven out of the eight testing windows, and by big margins. It is clear that after leakage is removed, not only both models’ performance decreases, but their comparative results also change to the opposite. In this case, the train-test leakage would have led to a completely different conclusion being drawn about the comparative advantage between the GCN and RF models on Android malware detection.
@@ 5.3 Impact of the Leakage on Performance Results
Window Model Architecture Measurement on Complete Test Data Performance (%) Measurement after Leak Removed from Test Performance (%)
@@
6 Leak-aware ML-based Android Malware Detector
In Section 1 we set out to answer two questions: 1) how will train-test leakage impact an ML model’s measured performance for Android malware detection, and 2) given that such leakage will happen “naturally” when the model is applied to real-world data, is there any practical value in discerning such differences of performance measurement? At this point we have clearly answered question 1): not only does the train-test leakage inflate measured performance, it may even flip the ultimate research conclusion that can be drawn. In this section we look at question 2). In our experiments we remove the leakage apps from testing data before evaluating a model’s performance. These leakage apps would be present in the real world data when the ML model is applied; wouldn’t evaluation on the complete test data (including the leakage apps) reflect the model’s utility in the real world? This question is especially releavant when research conclusion about different models’ comparative advantages flips when leakage data is removed (e.g., in case study 2). Should practitioners care about this result, if all they need is a model that would perform better on the real world data?
@@ 6 Leak-aware ML-based Android Malware Detector
In real-world operation, all data must be processed. Removing the leakage from the test set is intended to evaluate the ML model’s capacity to generalize – to determine whether good performance is due to the model’s ability to generalize to unseen samples or simply because it has memorized the data it has already seen. This differentiation is important for deciding which ML model to use, or whether to use ML at all. It will be helpful to understand why the comparative results flipped in case study 2. Figure 5 illustrates the performance of RF and GCN models on 1) the complete test data, 2) the leakage portion of the test data, and 3) the non-leakage portion of the test data. Non-leakage portion is the test data after the leakage is removed. The figure shows that both models achieve their highest performance when tested on the leakage portion of the test data and the lowest on the non-leakage portion of test data. But their comparative differences are opposite between the two portions. RF outperforms GCN on the leakage portion, indicating that RF is better than GCN at memorization. GCN outperforms RF on non-leakage portion, indicating that GCN is better than RF at generalization. It turns out that RF’s advantage at memorization outweights its disadvantage on generalization, and shows stronger overall performance on the complete test data. The question is, does this mean RF should be recommended as having better utility for real-world operation?
From a machine learning perspective, GCN is a better model since it can generalize better to data that is not highly similar to what it is trained on. When this model is applied to real-world data, inevitably some input app will be highly similar to some apps it is trained on (leakage). In such cases we could directly leverage the known labels of the highly similar apps in the training data to make a prediction, e.g., by taking a majority vote. This could perform even better than using the trained model for these leakage apps. In other word, GCN’s disadvantage in memorization may not matter, if we can find a different way to handle the apps that share identical or nearly identical representation to apps whose maliciousness is already known. Specifically, if an app’s representation is duplicate or highly similar to one from training, it could be more effective to directly use what we know about these highly similar training apps to make a prediction. Only when the app’s representation is not duplicate or similar to any app in the training, we use the ML model to leverage its ability to generalize.
To capitalize on this, we propose a leak-aware detector to improve overall detection performance. For an app the detector needs to make a prediction on, we first determine if the app is considered leakage, i.e., is there an app in the training data that has identical, or nearly identical representation to the app? If so we identify all such apps in the training data and use their labels’ majority vote to make a prediction. Otherwise we use the trained machine learning model to make a prediction. The intuition behind this approach is, for the apps that are train-test leakage, utilizing majority vote of the similar training apps may yield better performance than applying the machine learned model to predict these apps’ maliciousness. This could improve the overall performance of the leak-aware detector on the complete test data, including both leak and non-leak apps. To implement this leak-aware detector, the detecting system needs to constantly keep track of the set of apps that have been used in training the machine learning model, similar to what we did in our experiments.
@@
We implemented such leak-aware detectors for the models in case study 2, to answer the question of “whether the flipped research conclusion makes any difference in the real world.” For each model – RF and GCN, we utilize the same leakage definition – $Leak_{\textit{fv}}\cup Leak_{\textit{emb}}$ to build the leak-aware detector described above. The performance results of the two leak-aware detectors, along with the stand-alone versions (applying the ML model on all test data), are shown in Figure 6. The leak-aware GCN detector’s overall performance on the complete test data exceeds both the stand-alone RF and the leak-aware RF. This is in contrast to the comparison of stand-alone GCN and stand-alone RF, where RF outperforms GCN on the complete test data.
@@
This result shows that not only GCN outperforms RF on the testing data with leakage removed, it also outperforms RF on the complete test data when used in the leak-aware detector. Thus, the flipped comparative result between the two models is not only relevant to understanding the models’ ability to generalize to data it has not seen, it also has practical implications when the models are applied to real-world data that consist of both leakage and non-leakage apps. Specifically, by identifying the leakage apps, one can build a leak-aware detector that allows the GCN’s comparative strength against RF to show on the real-world data as well.
7 Why So Much Leakage?

### v2

@@ V Case Study 2: Graph Convolutional Networks for Android Malware Detection
V-A Summary of the Research and Dataset Used
@@
GCNs leverage graph structures as input, which allows for potentially capturing richer static code semantics, such as that expressed in an Android app’s ICFG. We use an Android app’s ICFG444 To obtain a single ICFG for an Android app, we combine the ICFG of each component of the app based on the life cycle semantics of how the methods of various components are called by the Android runtime. as the GCN model’s input, where each node represents an API call, with additional attributes including class path, method name, parameter type, and node degree. Following the standard GCN training process, we obtain a graph embedding for the ICFG, achieved through multiple rounds of message passing within the GCN, where each node’s embedding is updated based on its neighbors. Finally, a readout layer aggregates these node embeddings into a unified graph embedding for the entire ICFG. This component of the model can be seen as a graph encoder. The unified graph embedding learned by the encoder is then fed as input to the classification layer of the model. More details about the GCN model can be found in the Appendix B.
We compare the GCN approach to a standard ML approach – random forest (RF). The RF model takes as input a feature vector representing an app. We extract DREBIN-like static features and rank them in decreasing order of their mutual information with the class label. We then select 2,970 most informative features for training an RF model, which include permissions, intent actions, discriminative APIs, obfuscation signatures, and native code signatures. The mutual information is close to 0 for features ranked lower.
[…]
The complete dataset encompasses 24 batches, totaling 5,760 malicious and 90,240 benign apps, all chronologically ordered by their release date on Google Play. This dataset provides eight sliding windows (the window slides two batches each time) for evaluating machine learning models through different training, validation, and testing data. This sliding window approach, as the sampling approach for batch construction, also avoids the temporal bias [11] as discussed in Section II-B.
V-B Identification of Train-Test Leakage
@@
The two machine learning models our research compares, GCN and RF, use different input representations for an Android app. In GCN, an input app is represented as an ICFG adjacency matrix, together with a list of attributes for each node, whereas in RF an input app is represented as a feature vector. In prior discussion, we have explained the definition of train-test leakage when a feature vector is used as an app’s representation. These are apps in the testing data that share identical feature vectors with at least one app in the training data. We denote these leakage apps as $Leak_{\textit{fv}}$. While this definition works well for RF, it is not suitable to use in identifying train-test leakage for the GCN-based models since there an app is represented as a graph. Given that there is currently no known efficient algorithm to solve graph isomorphism, it is not clear how to efficiently determine if two Android apps’ ICFGs are identical. However, as mentioned above, the GCN model consists of an encoder and a classification layer. The encoder generates a graph embedding, i.e., a vector of floating numbers, which is provided as input to the classification layer.
Instead of comparing the ICFGs directly, we have chosen to use the computed graph embeddings of the ICFGs to identify train-test leakage. Specifically, given the embeddings $emb_{1}$ and $emb_{2}$ of two Android apps’ ICFGs, we compute cosine similarity between the two embeddings as
$\textit{cosine similarity}=\frac{emb_{1}\cdot emb_{2}}{\|emb_{1}\|\|emb_{2}\|}$
@@
where $\cdot$ is the inner product operation and $\|v\|$ represents the length of a vector $v$. The cosine similarity measures how similar the two vectors are and ranges from 0 to 111, where 111 means the two vectors are identical, while 0 means that the two vectors are orthogonal and share no similarity.
@@
Note that due to GCN model’s stocasticity and the numerical precision limitations, even if we feed the exact same ICFG multiple times to a GCN, the model will not generally produce the exact same graph embeddings. Thus, for the purpose of identifying train-test leakage, we chose a threshold $M$. A test app is considered a leakage if there is at least one training app whose cosine similarity with the testing app is at least $M$. We denote these apps as $Leak_{\textit{emb}}$. For a given testing set, we choose $M$ so that the two sets of leakage apps in the testing data, $Leak_{\textit{fv}}$ and $Leak_{\textit{emb}}$, are most aligned. The $M$ value varies from one testing window (Figure 3) to another, but they are all very close to 111 (e.g., 0.99980.99980.9998). More details on how the threshold $M$ is determined can be found in Appendix C.
@@
Figure 4 illustrates the ratio of leakage apps ($Leak_{\textit{fv}}$, $Leak_{\textit{emb}}$, and their union) on the testing data used in our experiments. A significant portion of apps in the testing data (oftentimes exceeding half of the total apps) are identified as train-test leakage. Since the two models give rise to two different sets of leakage apps, we need to remove the union of the two leakage sets from the testing data in order to compare the two models’ performance after leakage removal, for the same reasons as explained in Section IV-B.
V-C Impact of the Leakage on Performance Results
@@
Window
Model
Architecture
Measurement on Complete Test Data
Performance (%)
Measurement after Leak Removed from Test
Performance (%)
@@
Like in Section IV, we compare the performance measurements of the two models both before and after leakage apps are removed from testing data. Leakage portion is defined as $Leak_{\textit{fv}}\cup Leak_{\textit{emb}}$. Table III shows the results. For each testing window, the color-coded performance differences are included in parentheses, with green in favor of GCN and red in favor of RF. If we measure the performance on all the test data, RF consistently outperforms GCN based on the F1 score and balanced accuracy metric. However, once the leakage is removed, the comparative result flips: GCN outperforms RF in seven out of the eight testing windows, and by big margins. It is clear that after leakage is removed, not only both models’ performance decreases, but their comparative results also change to the opposite. In this case, the train-test leakage would have led to a completely different conclusion being drawn about the comparative advantage between the GCN and RF models on Android malware detection.
@@
VI Which is Important: Memorization or Generalization?
In Section I we set out to answer two questions: 1) how will train-test leakage impact an ML model’s measured performance for Android malware detection, and 2) given that such leakage will happen “naturally” when the model is applied to real-world data, is there any practical value in discerning such differences of performance measurement? At this point we have clearly answered question 1). In this section we look at question 2). In our experiments we remove the leakage apps from testing data before evaluating a model’s performance. These leakage apps would be present in the real world data when the ML model is applied; wouldn’t evaluation on the complete test data (including the leakage apps) reflect the model’s utility in the real world? This question is especially relevant when research conclusion about different models’ comparative advantages flips when leakage data is removed (e.g., in case study 2). Should practitioners care about this result, if all they need is a model that would perform better on the real world data?
In real-world operation, all data must be processed. Removing the leakage from the test set is intended to evaluate the ML model’s capacity to generalize – to determine whether good performance is due to the model’s ability to generalize to unseen samples or simply because it has memorized the data it has already seen. This differentiation is important for deciding which ML model to use, or whether to use ML at all. It will be helpful to understand why the comparative results flipped in case study 2. Figure 5 illustrates the performance of RF and GCN models on 1) the complete test data, 2) the leakage portion of the test data, and 3) the non-leakage portion of the test data. Non-leakage portion is the test data after the leakage is removed. The figure shows that both models achieve their highest performance when tested on the leakage portion of the test data and the lowest on the non-leakage portion of test data. But their comparative differences are opposite between the two portions. RF outperforms GCN on the leakage portion, indicating that RF is better than GCN at memorization. GCN outperforms RF on non-leakage portion, indicating that GCN is better than RF at generalization. It turns out that RF’s advantage at memorization outweights its disadvantage on generalization, and shows stronger overall performance on the complete test data. The question is, does this mean RF should be recommended as having better utility for real-world operation? In other words, which capability is more important for the practical utility of a machine learning model: memorization or generalization? The answer is resoundinlgy the latter. This is because for apps that are known malware, there are other more efficient methods to detect them, such as through a known signature. Only when a malware has not been widely recognized a machine learning model could potentially provide useful guidance. Thus, evaluating the machine learning model on the non-leakage part of the testing data is more meaningful than evaluating on the complete testing data.
To further illustrate this point, we create a hypothetical detector, which we call “leak-aware detector,” for the purpose of approximating the effect of combining machine learning and signature-based approaches to detect Android malware. For an app the detector needs to make a decision on, we first determine if the app is considered leakage, i.e., is there an app in the training data that has identical or highly similar representation to it? If so we identify all such apps and use their labels’ majority vote to make a decision. Otherwise we use the trained machine learning model to make a prediction. This clearly separates generalization and memorization: we only use machine learning for generalizing, and use signature matching for memorizing. This leak-aware scheme more accurately reflects a machine learning model’s utility in practice. It is designed as an approximate model for combining machine learning with signature detection and used as a measurement tool only. It is by no means an indication on how such combination shall be implemented for operations. In fact it is not an efficient implementation since the detector needs to constantly keep track of the set of apps that have been used in training the machine learning model. We implemented such leak-aware detectors for the models in case study 2, to answer the question of “whether the flipped research conclusion makes any difference in the real world.” For each model – RF and GCN, we utilize the same leakage definition – $Leak_{\textit{fv}}\cup Leak_{\textit{emb}}$ to build the leak-aware detector described above.
@@ VI Which is Important: Memorization or Generalization?
The performance results of the two leak-aware detectors, along with the stand-alone versions (applying the ML model on all test data), are shown in Figure 6. The leak-aware GCN detector’s overall performance on the complete test data exceeds both the stand-alone RF and the leak-aware RF. This is in contrast to the comparison of stand-alone GCN and stand-alone RF, where RF outperforms GCN on the complete test data.
@@
This result shows that not only GCN outperforms RF on the testing data with leakage removed, it also outperforms RF on the complete test data when used in the leak-aware scheme. Thus, the flipped comparative result between the two models is not only relevant to understanding the models’ ability to generalize to data it has not seen, it also has practical implications when the models are applied to real-world data and in combination with signature-based detection schemes. Specifically, by approximating the effect of such combination, the leak-aware detector allows the GCN’s comparative strength against RF to show on the complete testing data as well.
VII Related Work
VIII Conclusion
This work highlights a prevalent issue of train-test data leakage in machine learning (ML)-based Android malware detection. We demonstrate this issue through two case studies, illustrating that such leakage can inflate model performance and sometimes also lead to misleading research conclusions. The key issue is that the existence of train-test leakage causes the evaluation results to conflate a model’s ability to generalize with its ability to memorize. Thus it is more meaningful to evaluate a machine learning model’s performance on testing data with the leakage removed, than on the complete testing data. We further use a leak-aware detector scheme to approximate the effect of combining machine learning with signature-based detection, and shows more clearly which model can provide more practical utility in real-world use scenarios.
Appendix A: Evaluation Metrics Definition
First some common terminologies:
TP: True Positives (correctly classified malicious apps)
FN: False Negatives (malware mis-classified as benign)
TN: True Negatives (correctly classified benign apps)
FP: False Positives (benign apps mis-classified as malware)
$\textit{Precision}=\frac{\textit{TP}}{\textit{TP}+\textit{FP}}\hskip 14.45377pt\textit{Recall}=\frac{\textit{TP}}{\textit{TP}+\textit{FN}}$
[…]
Appendix B: Details about Our GCN Model
Our GCN model utilizes a four-layer graph convolutional architecture followed by a standard linear layer from the torch.nn library. ReLU (Rectified Linear Unit) activation ($max(x,0)$) was employed in the convolutional layers to extract localized node embeddings. We experimented with varying the number of convolutional layers (2-7) and hidden channels (32-128) within the model. While increasing hidden channels improved efficiency, GPU memory limitations prevented exploration beyond that. We found optimal performance with four convolutional layers and 64 hidden channels. Hyperparameter tuning identified 200 epochs and a learning rate of 0.001 as optimal for our dataset.
Appendix C: Embedding Similarity Threshold
In Section V, $M$ defines the level of similarity between graph embeddings for the purpose of considering train-test leakage. The following details the process of establishing this threshold for each data window.
For a given data window, an $M$ value will give rise to a subset of apps $\mathbf{Leak_{\textit{emb}}}$ in the testing data that are considered leakage for the GCN model. On the other hand, the subset $\mathbf{Leak_{\textit{fv}}}$ captures the leakage for the RF model. Intuitively, being leakage means that a testing app is very similar to a training app, whether under the RF model or the GCN model. Thus we’d like to choose an $M$ so that the corresponding subset $\mathbf{Leak_{\textit{emb}}}$ most closely matches the subset $\mathbf{Leak_{\textit{fv}}}$. We use the metric Intersection over Union (IoU) to measure how closely the two sets are aligned:
$\textit{Intersection over Union}\left(\textit{IoU}\right)=\frac{\mathbf{Leak_{\textit{fv}}}\cap\mathbf{Leak_{\textit{emb}}}}{\mathbf{Leak_{\textit{fv}}}\cup\mathbf{Leak_{\textit{emb}}}}$
IoU represents the proportion of elements that are common to $\mathbf{Leak_{\textit{fv}}}$ and $\mathbf{Leak_{\textit{emb}}}$ relative to their total combined size. It is a normalized score between 0 and 1. When the two sets completely coincide, IoU is 111. When the two sets are disjoint, IoU is 0.
We follow the steps below to find the threshold $M$ that maximizes IoU:
  1. 1.
    Define a similarity range (e.g., [0.8, 1.0]).
  2. 2.
    Set a small increment rate (e.g., 0.01) to step through the range.
  3. 3.
    At each step, calculate the IoU based on the current threshold value $M$.
  4. 4.
    We identify the $M$ that leads to the maximum IoU.
The above steps can be achieved through Algorithm 1. Figure 7 illustrates how we find thethreshold $M$ for a given experiment window.
Algorithm 1 Finding the Optimal Threshold $M$
1:Define a similarity range $R$ (e.g., [0.8, 1.0]).
2:Set a small increment rate r (e.g., 0.01) to step through the range.
3:Initialize $M_{\text{max}}$ to 0, $\text{IoU}_{\text{max}}$ to 0
4:for $M$ in range from $R_{\text{min}}$ to $R_{\text{max}}$ with step $r$ do
5:  Calculate IoU from current $M$ and $Leak_{\textit{emb}}$.
6:  if $\text{IoU}>\text{IoU}_{\text{max}}$ then
7:   $\text{IoU}_{\text{max}}$ = IoU
8:   $M_{\text{max}}$ = $M$
9:  end if
10:end for
11:Output: The optimal threshold $M_{\text{max}}$ that leads to the maximum IoU.
Figure 7: Determining the cosine similarity threshold ($M$) within the experiment window. $M$ is the value that leads to the maximum IoU.
Win_1 Win_2 Win_3 Win_4 Win_5 Win_6 Win_7 Win_8
Default Test 8,000 8,000 8,000 8,000 8,000 8,000 8,000 8,000
$\mathbf{Leak_{\textit{fv}}}$ 3,245 3,001 3,134 3,865 4,297 3,030 2,132 2,552
$\mathbf{Leak_{\textit{emb}}}$ 4,141 3,585 4,181 4,701 5,514 3,180 3,271 4,364
$\mathbf{Leak_{\textit{fv}}\cup Leak_{\textit{emb}}}$ 4,694 4,203 4,929 5,362 6,066 4,529 4,395 5,256
TABLE IV: Statistics of test data on each window. $\mathbf{Leak_{\textit{emb}}}$ is determined for each experiment window by selecting the $M$ value that maximizes the IoU score.
Train Test Win_1 Win_2 Win_3 Win_4 Win_5 Win_6 Win_7 Win_8
Win_1 0.4056 0.2699 0.1719 0.1596 0.1535 0.1301 0.0568 0.0321
Win_2 0.3751 0.2256 0.2116 0.1900 0.1585 0.0636 0.0405
Win_3 0.3917 0.3275 0.2670 0.1874 0.0811 0.0549
Win_4 0.4831 0.3698 0.2285 0.1140 0.0620
Win_5 0.5371 0.2850 0.1295 0.0770
Win_6 0.3787 0.1565 0.1003
Win_7 0.2665 0.1224
Win_8 0.3190
TABLE V: $Leak_{\textit{fv}}$ ratio in different test window while keep the training data fixed. A decreasing ratio of $Leak_{\textit{fv}}$ as the test data release timeline progresses.
Table IV summarizes the appearance of $\textit{Leak}_{\textit{fv}}$ or $\textit{Leak}_{\textit{emb}}$ in our test data. Table V shows a consistent decline in the $\textit{Leak}_{\textit{fv}}$ ratio within the test data over time, considering a fixed training dataset. For instance, training models solely on window-1 data leads to progressively lower $\textit{Leak}_{\textit{fv}}$ occurrences in newer test datasets. The $\textit{Leak}_{\textit{emb}}$ ratio follows a similar trend.
Appendix D: Why So Much Leakage?